Implement the boolean function F(x,y,z) = xy + x'z + yz

write all the steps
and identify the law rule
Answer #1
F(x,y,z) = xy + x'z + yz
Applying identity law
F(x,y,z) = xy + x'z + yz(1)
Applying inverse law
F(x,y,z) = xy + x'z + yz(x+x')
Applying distributive law
F(x,y,z) = xy + x'z + xyz + x'yz
Applying distributive law
F(x,y,z) = xy + x'z(1+y) + xyz
Applying identity law
F(x,y,z) = xy + x'z(1) + xyz
Applying identity law
F(x,y,z) = xy + x'z + xyz
Applying distributive law
F(x,y,z) = xy(1+z) + x'z
Applying identity law
F(x,y,z) = xy(1) + x'z
Applying identity law
F(x,y,z) = xy + x'z